# EmojiDiff: Advanced Facial Expression Control with High Identity Preservation in Portrait Generation

### v2

@@ 3.2.1 Base E-Adapter Training
In this subsection, we first propose the expression controller (E-Adapter) suitable for portrait generation and then train its basic version for subsequent data synthesis. Given the strong coupling between identity and expression, we employ a decoupled parallel attention to integrate them and design the E-Adapter, building upon IP-Adapter [49]. As illustrated in Fig. 3, E-Adapter includes two main control branches: the identity and expression branch. This simple yet effective design achieves three objectives. First, the parallel control of expression and identity reduces operation errors compared to the serial way. Second, the model can efficiently retrieve features from identity or expression, and learn how to incorporate the expression into the given identity. Third, the training module fully decouples expression and identity control, enabling independent or simultaneous manipulation of both (more empirical analysis can be found in Sec. 6). Specifically, in the identity branch, we inherited IP-Adapter to encode identity embeddings and inject them into the U-Net along with text embeddings, formulated as:
@@
Based on the designed controller, we utilize the same-identity data $\{\mathbf{S}_{k}^{g^{\prime}},\mathbf{R}_{k}^{g},\mathbf{D}_{k}^{g}\}$ to train the Base E-Adapter, adopting denoising loss $\mathcal{L}_{sd}$ throughout the process.
@@ 3.2.2 ID-irrelevant Data Iteration
The image identity constructed directly using the Base E-Adapter is greatly influenced by the expression reference’s appearance. The core idea of IDI is to minimize this disruption. To achieve this, we innovatively reuse Base E-Adapter to modify the identity of the generated image while keeping the expression unchanged. Specifically, we utilize expression images $\mathbf{R}_{k}^{g}$ inherited from training as the expression input but randomly selected portraits as the input of the identity branch, combining with the diffusion model to synthesize the restruct image. Subsequently, we perform FaceFusion [12] on the restruct image to create the faceswap image, further increasing the identity disparity between the generated image and the expression reference. Meanwhile, we employ facial blendshapes [16] and landmark [30] differences to exclude unqualified data (detailed in supplementary materia). Through these processes, we acquire cross-identity expression pairs $\{\mathbf{R}_{k^{\prime}}^{g},\mathbf{D}_{k}^{g}\}$ that maintain highly consistent expressions while exhibiting completely distinct identity, and the example is illustrated in Fig. 4.
Why can IDI keep the expression unchanged while changing the image’s identity? On one hand, as our insights at the beginning of Sec. 3.2, E-Adapter trained with the same identity data exhibits superior expression transfer capability, since the model tends to replicate the appearance and texture of the reference image. Additionally, the expression branch of the Base E-Adapter has been overfitted to all expression images by training. Consequently, IDI can effortlessly transmit the expression into the restruct images. On the other hand, the ID branch possesses certain identity control capabilities. The key is that the identity of the generated image does not need to match the input ID condition precisely; it simply needs to differ from that of the expression condition. By reusing the ID branch, we inject high-level semantic features of randomized portraits (e.g., gender, hairstyle, and face shape) into the restruct images. Moreover, the FaceFusion operation further alters the low-level facial textures, ensuring that the identity of the faceswap image is entirely distinct from the input.
Figure 4: Examples of IDI, detailed in Sec. 3.2.2.
@@ 3.3 ID-enhanced Contrast Alignment
To this end, we propose ICA to fine-tune the E-Adapter. Based on the same inputs, ICA incorporates the frozen Refined E-Adapter and the original identity controller (IP-Adapter [49]) as supervision to calculate expression loss $\mathcal{L}_{exp}$ and identity loss $\mathcal{L}_{id}$. $\mathcal{L}_{id}$ encourages the IP-Adapter and the E-Adapter which integrates expression control to generate similar portraits, thereby minimizing the impact of expression control on identity fidelity. Additionally, $\mathcal{L}_{exp}$ ensures that fine-tuning does not compromise the existing expression control capability of the Refined E-Adapter.
@@
$f(t)$ =$=\begin{cases}\sqrt{\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}}&\text{if }t\leq R_{tmax},\\ f(R_{tmax})&\text{if }t>R_{tmax},\end{cases}$
where ”${\rm Decode}$” is the VAE decode function, and $R_{tmax}$ is a constant. Through ANI, we can directly reconstruct images $\hat{x}_{0}$ at any given timestep $t$, even with large timesteps characterized by high noise. We present theoretical analysis of this mechanism in supplementary materia Sec. B.2.
@@
where $\cos$ is cosine similarity, and $\phi(\cdot)$ is the pre-trained face embedding extractor [10]. Similarly, we utilize the frozen Refined E-Adapter to generate image $\hat{x}_{0}^{exp}$. The expression loss $\mathcal{L}_{exp}$ reduces the difference in facial landmarks between $\hat{x}_{0}^{exp}$ and $\hat{x}_{0}^{cur}$, is defined as follows:
@@
where ”${\rm Wing}$” denotes the wing loss [13], and $\varphi(\cdot)$ is the pre-trained facial landmarks detector [42].
@@
MediaPipe [30] 0.623 4.985 0.104 0.472 0.087 4.713 0.149 0.671 0.287 4.476 0.130 0.524
Arc2Face [10] 0.618 4.946 0.116 0.537 0.097 4.630 0.163 0.713 0.146 4.349 0.141 0.609
FineFace [43] 0.532 4.870 0.120 0.576 0.090 4.423 0.156 0.700 0.297 3.813 0.142 0.624
@@ 4.1 Experiment Setting
Dataset. We first collect 10k indoor and outdoor images of people of different countries and genders, each with 3-12 expressions. Then, we use LIQE [52] to filter out low-quality images and crop out 512$\times$512 face square images. Subsequently, this dataset is employed to train the Base E-Adapter based on SD1.5 [38]. Afterward, we construct about 120k cross-identity triplets through IDI. Through expression consistency filtering, we retain about 100k triplets named CIEP100K, for training and fine-tuning the Refined E-Adapter. For evaluation, we gather images of 50 celebrities from diverse fields, with 30 different expressions.
@@
Implementation Details. We train the Base E-Adapter on SD1.5 to construct data and the Refined E-Adapter on both SD1.5 [38] and SDXL [36]. In the Base E-Adapter training stage, our model is trained from scratch for 0.5 days. In SD1.5, we train the Refined E-Adapter for 1 day and fine-tune it for 4 hours. In SDXL, the input images are super-resolution using SwinIR [27] to a resolution of 1024$\times$1024, and we train the Refined E-Adapter for 2 days and fine-tune it for 0.5 days. More details are provided in supplementary materia Sec. B.1.
@@
Baselines. We compare our model with prior methods, including OpenPose [51], MediaPipe [30], Arc2Face [33], and FineFace [43], using the same configurations (base model, text prompt, inference steps, etc.). Although portrait animation methods are fundamentally different from our approach, we assess these methods, including FaceAdapter [19], X-Portrait [47], and LivePortrait [17], to thoroughly evaluate our method. For the diffusion-based methods, we integrate IP-Adapter [49] plugin for identity preservation. For animation-based methods, an expressionless image of the portrait is generated, and then animation is applied. We select LivePortrait [17] to construct data triples $\mathcal{T}$ and train the E-Adapter, which serves as our strong baseline.
@@
Benchmarks. To ensure a fair comparison, all experiments are conducted using the SD1.5 framework across three styles: realistic, anime, and ink painting, while the style is directly dependent on the base model and LoRA. Given the identity image, expression reference, and the generated image, we evaluate all methods using ID fidelity (ID), image quality (IQ), expression similarity (Exp.), and facial landmark movements similarity (LMS). Different from the models used in the ID encoder, the Antelopev2 [1] model is utilized to extract identity embeddings from both the portrait and generated image, with the cosine similarity between them used to evaluate identity fidelity. For the assessment of image quality, the pre-trained network LIQE [52] is employed. Expression similarity is measured by computing the L1 difference between facial blendshapes extracted from the expression and generated image using MediaPipe [30]. To evaluate the keypoint similarity, crucial facial landmarks (such as eyes, pupils, and mouth) are detected with MediaPipe to calculate movement amplitude differences. More metric details are provided in supplementary Sec. A.
@@ 4.2 Comparison
Quantitative results. As summarized in Tab. 1, our method consistently outperforms all competitors by a good margin. Compared to the same type of diffusion conditioning methods, our approach shows significant improvements in ID similarity, image quality, and expression control ability. Even evaluating against portrait animation methods such as FaceAdapter [19] and LivePortrait [17], the proposed method offers superior ID fidelity and more accurate expression control. Notably, X-Portrait, which directly uses RGB expression images, also demonstrates strong expression control. However, X-Portrait shows a noticeable drop in image quality when applied to non-realistic styles, with IQ scores of only 4.142 and 3.668 in the anime and painting styles, respectively. On the contrary, our method produces exquisite stylized images due to the seamless integration with the diffusion model.
Qualitative results. Fig. 5 illustrates the performance of various methods in ID and expression control across three styles. These examples demonstrate that our model effectively conveys reference expressions, including lip movements and eye gaze while preserving the identity information of the source portrait. The fourth and sixth examples in the figure highlight the ability of our method to capture subtle movements, including lower lip pursing and pouting. Compared to other approaches, our model consistently exhibits robust expression control across different styles.
@@ 4.3 Ablation Study
We evaluate the proposed IDI and ICA on SD1.5 and SDXL. As presented in Tab. 2, in comparison to our baseline, IDI substantially improves expression control, reducing the Exp. score of the two base models by 23.0% and 22.6%, respectively. Additionally, ICA further enhances ID fidelity without compromising expression control, increasing the corresponding ID score by 3% and 6%, respectively. Similarly, as illustrated in Fig. 8, the complete method retains more identity details, including glasses and hair color, compared to the method without ICA.
IDI ICA SD1.5 [38] SDXL [36]
@@
Figure 6: Different expression controller structure.
ID-irrelevant Data Iteration. As shown in Tab. 3, we assess the performance of E-Adapter at different data iteration stages. Initially, the Base E-Adapter achieves significant reductions in Exp. and LMS scores compared to the baseline, showing the powerful expression transfer ability of Base E-Adapter. Additionally, by randomizing the ID branch’s input for cross-identity data construction, the newly trained Refined E-Adapter substantially reduces ID leakage and improves the ID score by 0.136. Finally, through further face-swapping, the Refined E-Adapter significantly outperforms the baseline.
Expression Controller Architecture. As illustrated in Fig. 6, we experiment with various expression controller structures. Expression embeddings are injected into the diffusion model using one of the following strategies: (a) self-attention, (b) ControlNet, (c) serial cross-attention with ID embeddings, and (d) parallel cross-attention with ID embeddings. As indicated in Tab. 4, the schemes (a) and (c), which operate without parallel integration with ID embeddings, are insufficient in both ID fidelity and expression control capability. While the ControlNet achieves high ID fidelity, it lacks satisfactory expression control. Our approach combines identity and expression embeddings within a unified cross-attention block, effectively fostering their interaction and delivering optimal performance.
Generalizability of Expression Control. As depicted in Fig. 7, prompts can be customized to meet diverse requirements while expressions are stably governed by the expression template. Notably, an appropriate prompt will further enhance the vividness of expressions, and a conflicting prompt only slightly affects expressions, as demonstrated in the last two columns of Fig. 7. Moreover, the third and fourth columns of Fig. 8 demonstrate that the identity and expression branch can independently perform their respective functions, enhancing the flexibility of facial control.
Base
E-Aadapter
Restruct
Image
FaceSwap
Image
ID $\uparrow$ Exp. $\downarrow$ LMS $\downarrow$
0.265 0.113 0.447
$\surd$ 0.127 0.065 0.283
$\surd$ $\surd$ 0.263 0.080 0.362
$\surd$ $\surd$ $\surd$ 0.295 0.087 0.355
@@
Figure 7: Images generated using different prompts.

### v3

@@ 3.2.1 Base E-Adapter Training
In this subsection, we first propose the expression controller (E-Adapter) suitable for portrait generation and then train its basic version for subsequent data synthesis. Given the strong coupling between identity and expression, we employ a decoupled parallel attention to integrate them and design the E-Adapter, building upon IP-Adapter [ye2023ipa]. As illustrated in Fig. 3, apart from text input, E-Adapter includes two control branches: the identity and expression branch. This simple yet effective design achieves three objectives. First, the parallel control of expression and identity reduces operation errors compared to the serial way. Second, the model can efficiently retrieve features from identity or expression, and learn how to incorporate the expression into the given identity. Third, the training module fully decouples expression and identity control, enabling independent or simultaneous manipulation of both (more empirical analysis can be found in Sec. 2 and Sec. 4.4). Specifically, in the identity branch, we inherited IP-Adapter to encode identity embeddings and inject them into the U-Net along with text embeddings, formulated as:
@@
Based on the designed controller, we utilize the same-identity data $\{\mathbf{S}_{i}^{\neg{e}},\mathbf{R}_{i}^{e},\mathbf{T}_{i}^{e}\}$ to train the Base E-Adapter, adopting denoising loss $\mathcal{L}_{sd}$ throughout the process.
@@ 3.2.2 ID-irrelevant Data Iteration
The image identity constructed using the Base E-Adapter is greatly influenced by the expression reference’s appearance. The core idea of IDI is to minimize this disruption. To achieve this, we innovatively reuse Base E-Adapter to modify the identity of the generated image while keeping the expression unchanged. Specifically, we utilize expression images $\mathbf{R}_{i}^{e}$ inherited from training as the expression input but random identity $\neg{i}$ as the input of the identity branch, combining with the diffusion model to synthesize the restruct image. Subsequently, we perform FaceFusion [FaceFusion] on the restruct image to create the faceswap image, further increasing the identity disparity between the generated image and the expression reference. Meanwhile, we employ facial blendshapes [BlendshapeV2] and landmark [lugaresi2019mediapipe] differences to exclude unqualified data (detailed in supplementary material Sec. A.1). Through these processes, we acquire cross-identity expression pairs $\{\mathbf{R}_{\neg{i}}^{e},\mathbf{T}_{i}^{e}\}$ that maintain highly consistent expressions while exhibiting completely distinct identity, and the example is illustrated in Fig. 4.
Why can IDI keep the expression unchanged while changing the image’s identity? On one hand, as our insights at the beginning of Sec. 3.2, E-Adapter trained with the same identity data exhibits superior expression transfer capability, since the model tends to replicate the appearance and texture of the reference image. Additionally, the expression branch of the Base E-Adapter has been overfit to all expression images by training. Consequently, IDI can effortlessly transmit the expression into the restruct images. On the other hand, the ID branch possesses certain identity control capabilities. By reusing the ID branch, we inject high-level semantic features of randomized portraits (e.g., gender, hairstyle, and face shape) into the restruct images. Notably, the identity of the generated image does not need to match the input ID condition precisely; it simply needs to differ from that of the expression condition. Moreover, the FaceFusion operation further alters the low-level facial textures, ensuring that the identity of the faceswap image is entirely distinct from the input.
Figure 4: Examples of IDI, detailed in Sec. 3.2.2
@@ 3.3 ID-enhanced Contrast Alignment
To this end, we propose ICA to fine-tune the E-Adapter. Based on the same inputs, ICA incorporates the frozen Refined E-Adapter and the original identity controller (IP-Adapter [ye2023ipa]) as supervision to calculate expression loss $\mathcal{L}_{exp}$ and identity loss $\mathcal{L}_{id}$. $\mathcal{L}_{id}$ encourages the IP-Adapter and the E-Adapter which integrates expression control to generate similar portraits, thereby minimizing the impact of expression control on identity fidelity. Additionally, $\mathcal{L}_{exp}$ ensures that fine-tuning does not compromise the existing expression control capability of the Refined E-Adapter.
@@
[…]
where $\rm Decode$}” is the VAE decode function, and $R_{tmax}$ is a constant. Through ANI, we can directly reconstruct images $\hat{x}_{0}$ at any given timestep $t$, even with large timesteps characterized by high noise. We present theoretical analysis of this mechanism in supplementary material Sec. B.
@@
where $\cos$ is cosine similarity, and $\phi(\cdot)$ is the pre-trained face embedding extractor [deng2019arcface]. Here, we use $\hat{x}_{0}^{id}$ rather than another clean portrait image as ground truth of $\mathcal{L}_{id}$, since we find that the noise discrepancy between the real image and one-step decoded image $\phi(\hat{x}_{0}^{cur})$ destabilizes fine-tuning (ablated in Sec. 4.3). Similarly, we utilize the frozen Refined E-Adapter to generate image $\hat{x}_{0}^{exp}$. The expression loss $\mathcal{L}_{exp}$ reduces the difference in facial landmarks between $\hat{x}_{0}^{exp}$ and $\hat{x}_{0}^{cur}$, is defined as follows:
@@
where $\rm Wing$}” denotes the wing loss [feng2018wing], and $\varphi(\cdot)$ is the pre-trained facial landmarks detector [song2022gratis].
@@
MediaPipe [lugaresi2019mediapipe] 0.623 4.985 0.104 0.472 0.087 4.713 0.149 0.671 0.287 4.476 0.130 0.524
Arc2Face [deng2019arcface] 0.618 4.946 0.116 0.537 0.097 4.630 0.163 0.713 0.146 4.349 0.141 0.609
FineFace [varanka2024fineface] 0.532 4.870 0.120 0.576 0.090 4.423 0.156 0.700 0.297 3.813 0.142 0.624
@@ 4.1 Experiment Setting
Dataset. We first collect 10k indoor and outdoor images of people of different countries and genders, each with 3 to 12 expressions. Then, we use LIQE [zhang2023liqe] to filter out low-quality images and crop out $\times512$ face square images. Subsequently, this dataset is employed to train the Base E-Adapter based on SD1.5 [rombach2022sd]. Afterward, we construct about 120k cross-identity triplets through IDI. Through expression consistency filtering, we retain about 100k triplets named CIEP100k, for training and fine-tuning the Refined E-Adapter. For evaluation, we gather images of 50 celebrities from diverse fields, with 30 different expressions.
@@
Implementation Details. We train the Base E-Adapter on SD1.5 to construct data and the Refined E-Adapter on both SD1.5 [rombach2022sd] and SDXL [podell2023sdxl]. In the Base E-Adapter training stage, our model is trained from scratch for 0.5 days. In SD1.5, we train the Refined E-Adapter for 1 day and fine-tune it for 4 hours. In SDXL, we train the Refined E-Adapter for 2 days and fine-tune it for 0.5 days. More details are provided in supplementary material Sec. A.1.
@@
Baselines. We compare our model with prior methods, including OpenPose [zhang2023controlnet], MediaPipe [lugaresi2019mediapipe], Arc2Face [paraperas2024arc2face], and FineFace [varanka2024fineface], using the same configurations (base model, text prompt, inference steps, etc.). Although portrait animation methods are fundamentally different from our approach, we assess these methods, including FaceAdapter [han2024faceadapter], X-Portrait [xie2024xportrait], and LivePortrait [guo2024liveportrait], to thoroughly evaluate our method. For the diffusion-based methods, we integrate IP-Adapter [ye2023ipa] plugin for identity preservation. For animation-based methods, an expressionless image of the portrait is generated, and then animation is applied. We select LivePortrait [guo2024liveportrait] to construct data triples $\mathcal{T}$ and train the E-Adapter, which serves as our strong baseline.
@@
Benchmarks. For a fair comparison, all experiments are conducted using the SD1.5 framework across three styles: realistic, anime, and ink painting, while the style is directly dependent on the base model and style LoRA. Given the identity image, expression reference, and the generated image, we evaluate all methods using ID fidelity (ID), image quality (IQ), expression similarity (Exp.), and facial landmark movements similarity (LMS). Different from ID encoder in E-Adapter, the Antelopev2 [Antelopev2] model is utilized to extract identity embeddings from both the portrait and generated image, with the cosine similarity between them used to evaluate identity fidelity. For the assessment of image quality, the pre-trained network LIQE [zhang2023liqe] is employed. Expression similarity is measured by computing the L1 difference between facial blendshapes extracted from the expression and the generated image using MediaPipe [lugaresi2019mediapipe]. To evaluate the keypoint similarity, crucial facial landmarks (such as eyes, pupils, and mouth) are detected [lugaresi2019mediapipe] to calculate movement amplitude differences. More metric details are provided in supplementary Sec. A.3.
@@ 4.2 Comparison
Quantitative results. As summarized in Tab. 1, our method consistently outperforms all competitors by a good margin. Compared to the same type of diffusion conditioning methods, our approach shows significant improvements in ID similarity, image quality, and expression control ability. Even evaluating against portrait animation methods such as FaceAdapter [han2024faceadapter] and LivePortrait [guo2024liveportrait], the proposed method offers superior ID fidelity and more accurate expression control. Notably, X-Portrait, which directly uses RGB expression images, also demonstrates strong expression control. However, X-Portrait shows a noticeable drop in image quality when applied to non-realistic styles, with IQ scores of only 4.142 and 3.668 in the anime and painting styles, respectively. On the contrary, our method produces exquisite stylized images due to the seamless integration with the diffusion model. We provide comparisons with more methods in the supplementary material Sec. C.1.
Qualitative results. Fig. 5 illustrates the results of various methods across three styles. These examples demonstrate that our model effectively conveys reference expressions, including lip movements and eye gaze while preserving the portrait identity. The fourth and sixth examples in the figure highlight the ability of our method to capture subtle movements, including lower lip pursing and pouting. Similarly, as illustrated in Fig. 6, the complete method retains more identity details(e.g., glasses and hair color), compared to the method without ICA.
@@ 4.3 Ablation Study
We evaluate the proposed IDI and ICA on SD1.5 and SDXL. As presented in Tab. 2, in comparison to our baseline, IDI substantially improves expression control, reducing the Exp. score of the two base models by 23.0% and 22.6%, respectively. Additionally, ICA further enhances ID fidelity without compromising expression control, increasing the corresponding ID score by 3% and 6%, respectively.
IDI ICA SD1.5 [rombach2022sd] SDXL [podell2023sdxl]
@@
ID-irrelevant Data Iteration. As shown in Tab. 3, we assess the performance of E-Adapter at different data iteration stages. Initially, the Base E-Adapter achieves significant reductions in Exp. and LMS scores compared to the baseline, showing the powerful expression transfer ability of Base E-Adapter. After using the constructed restruct images, the newly trained Refined E-Adapter substantially reduces ID leakage and improves the ID score by 0.136. Finally, through further face-swapping on the data, the Refined E-Adapter significantly outperforms the baseline.
ID-enhanced Contrast Alignment. As shown in Tab. 4, when utilizing clean real images as supervision, the ANI effectively improves ID fidelity and expression controllability. Additionally, using model predictions $\hat{x}_{0}^{id}$ rather than real images ensures consistency of noise levels between data pairs, further facilitating model fine-tuning.
Expression Controller Architecture. We experiment with various expression controller structures (The detailed structure can be found in the supplementary material Sec. A.2.). As indicated in Tab. 5, our approach combines identity and expression embeddings within a unified cross-attention block, effectively fostering their interaction and delivering optimal performance.
4.4 Generalizability of Proposed Approach
Decoupled identity and expression. As shown in Fig. 6, the third and fourth columns demonstrate that the identity and expression branch can take effect independently, enhancing the flexibility of facial control.
E-Adapter as the expression extractor. The expression embedding from our method can be generalized to expression-related tasks. Specifically, we evaluate the expression module of the E-Adapter on the expression recognition task following the classic method DLN [Zhang2021Learning]. Our method exhibits powerful expression recognition capabilities (88.7 vs DLN’s 86.4 in the RAF-DB benchmark [RAF]), demonstrating its efficacy as a robust expression extractor.
Generalization to other architecture. In the main experiment, considering generalizability of method and fair comparison with the SD1.5-based animation methods (e.g., XPortrait [xie2024xportrait], FaceAdapter [han2024faceadapter]), we simply injected condition information via the attention mechanism, which is similar to IP-Adapter [ye2023ipa]. As shown in Tab. 6, by directly utilizing the CIEP100k dataset and ICA, our method is seamlessly adopted to InstantID [wang2024instantid] and achieves substantially more detailed expression control.
Stage / Data Source ID $\uparrow$ Exp. $\downarrow$ LMS $\downarrow$
Baseline 0.265 0.113 0.447
Base E-Aadapter 0.127 0.065 0.283
using Restruct Image 0.263 0.080 0.362
Ours (using Faceswap Image) 0.295 0.087 0.355
@@ 4.4 Generalizability of Proposed Approach
Supervision ID $\uparrow$ Exp. $\downarrow$ LMS $\downarrow$
Clean Image 0.477 0.088 0.376
Clean Image w/ ANI 0.485 0.078 0.361
Ours 0.509 0.071 0.341
Table 4: Ablation study of ICA supervision.
@@
Method ID $\uparrow$ Exp. $\downarrow$ LMS $\downarrow$
PuLID [guo2024pulid] 0.536 0.117 0.502
PhotoMaker [li2024photomaker] 0.467 0.136 0.618
InstantID [wang2024instantid] 0.543 0.105 0.476
Ours 0.509 0.071 0.341
Ours w/ InstantID 0.540 0.077 0.356
Table 6: Comparisons with portrait generation methods.